What is the slope of a line perpendicular to the line 4x - 5y = 8?
Answer by checkley75(3666)

4X-5Y=8
-5Y=-4X+8
Y=-4X/-5+8/-5
Y=4X/5-8/5 THIS LINE HAS A SLOPE OF 4/5 THUS THE PERPENDICULAR LINE HAS A SLOPE OF -5/4 [THE NEGATIVE INVERSE].